# Length-Controlled AlpacaEval: A Simple Way to Debias Automatic Evaluators

### v1

@@ 4 Results
We apply our approach to AlpacaEval, as this benchmark has known length confounders, contains a large set of pre-computed LLM-based pairwise comparisons, and is widely used by the research community. We evaluate our approach on several measures of interest:
  • Decreasing length gameability: We call a metric length gameable if simply prompting the model to be more or less verbose significantly affects the metric outcome. Ideally, length gameability is low for two reasons. First, we would like evaluations that prioritize the content rather than the style of the answer. Second, the benchmark should not be too dependent on the prompting strategy as users usually think of evaluations of models rather than the entire system, which includes the prompt.
@@
  • Robustness and interpretability: Our corrected metric should be robust to simple adversarial attacks such as truncation, and be interpretable to users as a win-rate.
@@ 4.1 AlpacaEval-LC decreases length gameability
Figure 4: Closed source models often perform better (red) on length-controlled AlpacaEval as they are often shorter. The first column shows the length of outputs. The 4th and 5th columns, respectively, show the win rates and rank gains due to LC. The table shows the top 15 and bottom 5 systems in the leaderboard.
@@ 4.3 AlpacaEval-LC is interpretable and robust
Figure 5: With our GLM we can predict what the win rate would be if the baseline was any other model from the leaderboard.
Regularization makes LCAE robust to truncation
One potential issue with simple bias corrections is that they may be gamed through white-box adversarial attacks, e.g., postprocessing the outputs of models to make them look better on AlpacaEval-LC. One example of such an attack is to truncate all outputs to a few characters, besides those that are much better and around the same length as the baseline. A naive GLM fitted on such outputs should naturally predict very high win rates in the counterfactual world where outputs have the same length as the baseline. Indeed, when doing such post-processing to GPT-4 outputs, win rates increase from $3.7$ (AlpacaEval 2.0) to $25.9$ (AlpacaEval-LC, no regularization). To mitigate such adversarial attacks, our approach includes a regularization term on $\phi_{m,b}$. This decreases the gamed win rate to $12.2$ (AlpacaEval-LC with regularization) while having an imperceptible impact on standard models.
Interpretability as a win-rate
@@
4.4 Comparisons to baselines for length control
Figure 6: Length-controlled win rate has the best Arena Correlation and gameability from considered methods, while still being relatively robust to adversarial attacks.
Let’s now briefly discuss two other potential family length-correction methods that have been proposed in the community Duong (2024); Galambosi (2024); Teortaxes (2024).
@@ 4.4 Comparisons to baselines for length control
Another common way to control some covariates is through stratification. One potential metric, dubbed length-balanced (LB) win rate, would thus be to compute the average win rate stratified on examples where the model outputs are (1) longer and (2) shorter than the baseline Duong (2024). LB satisfies many of the desiderata of length control but has one main downside: robustness.
@@
The first and last rows in fig. 6 show that length-balanced win rates improve both the length gameability (measured by the normalized standard deviation of win rate across concise/standard/verbose prompts) and the Chatbot arena correlation. However, this approach is strictly dominated by our length-controlled method – in arena correlation, gameability, and adversarial win rate gains from truncating bad GPT-4 outputs as discussed in section 4.3.
@@
Figure 6 shows that this simple LN win rate performs surprisingly well on many of the metrics. We chose to present and implement the length-controlled (LC) win rate, as it is more principled (as an estimate of the direct effect), interpretable (as a win rate), and performs slightly better on all quantitative metrics except adversarial gameability.
@@ 5 Discussion
Figure 7: Length-controlled win rate has the best Arena Correlation and gameability from considered methods, while still being relatively robust to adversarial attacks.
@@
Length is a well-known bias of automated evaluators of chatbot LLMs but several others have been noted, including a bias of models towards their own outputs Zheng et al. (2023), or presence of lists Dubois et al. (2023). While we focus on a more detailed study of length biases here, we note that the same approaches can be applied to other biases by representing them as additional features in the logistic regression.
Additionally, our preliminary explorations of self-annotator biases shows that the effect exists but is often smaller than general model differences. fig. 7 shows that the ranking of considered models does not change when using different annotators. In particular, Claude 3 Opus prefers GPT4 Preview, and Mistral Large prefers the former two than itself.
@@
Our work is closely related to the recent work that aims to debias (implicit or explicit) reward models used to finetune LLMs with RLHF (Singhal et al. (2023)). For example Shen et al. (2023); Chen et al. (2024) try to train a reward model that is uncorrelated to length by making it predict the length at the same time as the reward and disentangle the two. Park et al. (2024) extends this intuition to the case of implicit reward models. This type of debiasing would not work out-of-the-box in typical auto-evaluation settings, e.g. AlpacaEval, which uses closed source LLM as judges rather than training a reward model. Our post-hoc debiasing could however be used in the RLHF setting, and we encourage future work to look into that.
@@ Acknowledgments
We thank OpenAI and Together AI for API credits to generate outputs and evaluate models. We thank the community for all the 100+ models they added to AlpacaEval. We thank Xuechen Li, Rohan Taori, and Tianyi Zhang for help maintaining AlpacaEval. We thank Viet Hoang Tran Duong for suggesting to consider length-balanced win rates. We thank the Twitter ML community for emphasizing the need of length-controlled autoevaluations. YD is supported by a Knights-Hennessy Scholarship.

### v2

@@ 4 Results
We apply our approach to all of AlpacaEval, as this benchmark has known length confounders, contains a large set of pre-computed LLM-based pairwise comparisons ($>120$ models, 805 instructions), and is widely used by the research community. We evaluate our approach on several measures of interest:
  • Decreasing length gameability: We call a metric length gameable if simply prompting the model to be more or less verbose significantly affects the metric outcome. Ideally, length gameability should be low for two reasons. First, we would like evaluations that prioritize the content rather than the style of the answer. Second, the benchmark should not be too dependent on the prompting strategy as users usually think of evaluations of models rather than the entire system, which includes the prompt.
@@
  • Robustness and interpretability: Our corrected metric should be robust to simple adversarial attacks such as truncation, and be interpretable to users as a win rate.
@@ 4.1 AlpacaEval-LC decreases length gameability
Figure 4: Closed source models often perform better (red) on length-controlled AlpacaEval as they are often shorter. The first column shows the length of outputs. The 4th and 5th columns, respectively, show the win rates and rank gains due to LC. The table shows the top 15 and bottom 5 systems (separated by the dark gray line) on AlpacaEval’s leaderboard from March 19th 2024.
@@ 4.3 AlpacaEval-LC is interpretable and robust
Figure 5: With our GLM we can predict what the win rate would be if the baseline was any other model from the leaderboard. The $x$-axis shows the baseline, and the $y$-axis the model. Both are in the same order.
Regularization makes AlpacaEval-LC robust to truncation
One potential issue with simple bias corrections is that they may be gamed through white-box adversarial attacks, e.g., postprocessing the outputs of models to make them look better on AlpacaEval-LC. One example of such an attack is to truncate all outputs to a few characters, besides those that are much better and around the same length as the baseline. A naive GLM fitted on such outputs should naturally predict very high win rates in the counterfactual world where outputs have the same length as the baseline. Indeed, when doing such post-processing to GPT-4 outputs, win rates increase from $3.7$ (AlpacaEval 2.0) to $25.9$ (AlpacaEval-LC, no regularization). To mitigate such adversarial attacks, our approach includes a regularization term on $\phi_{m,b}$. This decreases the gamed win rate to $12.2$ (AlpacaEval-LC with regularization) while having an imperceptible impact on non-adversarial models.
Interpretability as a win rate
@@
4.4 Different length control methods
Table 1: Length-controlled win rate has the best Chatbot Arena correlation and gameability from considered methods, while still being relatively robust to adversarial attacks. Each row is a different metric we considered.
  Chatbot Arena correlation ($\uparrow$) Gameability ($\downarrow$) Adversarial win rate gain ($\downarrow$)
 Win rate 0.94 26% 0.0
    Length-controlled 0.98 10% 8.5
    Length-normalized 0.96 15% 3.6
    Length-balanced 0.95 15% 40.8
Let’s now briefly discuss two other potential family length-correction methods that have been proposed in the community (Duong, 2024; Galambosi, 2024; Teortaxes, 2024).
@@ 4.4 Different length control methods
Another common way to control some covariates is through stratification. One potential metric, dubbed length-balanced (LB) win rate, would thus be to compute the average win rate stratified on examples where the model outputs are (1) longer and (2) shorter than the baseline (Duong, 2024). LB satisfies many of the desiderata of length control but has one main downside: robustness.
@@
The first and last rows in table 1 show that length-balanced win rates improve both the length gameability (measured by the normalized standard deviation of win rate across concise/standard/verbose prompts) and the Chatbot arena correlation. However, this approach is strictly dominated by our length-controlled method—in terms of correlation with Chatbot Arena, gameability, and adversarial win rate gains from truncating bad GPT-4 outputs as discussed in section 4.3.
@@
Table 1 shows that this simple LN win rate performs surprisingly well on many of the metrics. We chose to present and implement the length-controlled (LC) win rate, as it is more principled (as an estimate of the direct effect), interpretable (as a win rate), and performs slightly better on all quantitative metrics except adversarial gameability.
@@ 5 Discussion
Table 2: LLM judges self-bias but the ranking is surprisingly stable. Each column shows the leaderboard on AlpacaEval dataset as evaluated by different LLM judges. Rows show the win rate of the evaluated models against the judge.
  gpt-4-1106 claude-3-opus mistral-large
 gpt-4-1106 50.0 50.0 50.0
claude-3-opus 40.4 43.3 47.5
mistral-large 32.7 28.2 45.5
gpt-4-0613 30.2 20.5 34.3
gpt-3.5-turbo 19.3 16.7 28.9
@@
Length is a well-known bias of automated evaluators of chatbot LLMs but several others have been noted, including a bias of models towards their own outputs (Zheng et al., 2023), or presence of lists (Dubois et al., 2023). While we focus on a more detailed study of length biases here, we note that the same approaches can be applied to other biases by representing them as additional features in the logistic regression.
Additionally, our preliminary explorations of self-annotator biases show that the effect exists but is often smaller than general model differences. table 2 shows that the ranking of considered models does not change when using different annotators. In particular, Claude 3 Opus prefers GPT4 Preview, and Mistral Large prefers the former two than itself.
@@
Our work is closely related to the recent work that aims to debias (implicit or explicit) reward models used to finetune LLMs with RLHF (Singhal et al. (2023)). For example Shen et al. (2023); Chen et al. (2024) try to train a reward model that is uncorrelated to length by making it predict the length at the same time as the reward and disentangle the two. Park et al. (2024) extends this intuition to the case of implicit reward models. This type of debiasing would not work out-of-the-box in typical auto-evaluation settings, e.g. AlpacaEval, which uses closed-source LLM as judges rather than training a reward model. Our post-hoc debiasing could however be used in the RLHF setting, and we encourage future work to look into that.
Limitations
Firstly, we only tested our proposed debiasing mechanism on the AlapcaEval benchmark, which uses a set of relatively simple English instructions and a particular prompt for the LLM judge. Secondly, AlpacaEval-LC is based on the simplifying assumption that you would like to compare the model and the baseline as if they had the same length. Finally, we do not (aim to) solve any other issues associated with the use of an LLM-judge Zheng et al. (2023). Despite these limitations, we show that the correlation with Chatbot Arena increases significantly, which suggests that AlpacaEval-LC takes a step in the right direction.
@@ Acknowledgments
We thank Xuechen Li, Rohan Taori, and Tianyi Zhang for help maintaining AlpacaEval. We thank Viet Hoang Tran Duong for suggesting to consider length-balanced win rates. We thank the Twitter ML community for emphasizing the need of length-controlled autoevaluations. We thank the community for all the 100+ models they added to AlpacaEval. We thank OpenAI and Together AI for API credits to generate outputs and evaluate models. We gratefully acknowledge the support of an Open Philanthropy Project Award. Tatsunori Hashimoto is supported by a gift from Open Philanthropy and by the Tianqiao and Chrissy Chen Institute. Yann Dubois is supported by a Knights-Hennessy Scholarship.